I need help finding the discriminant of this problem:
-10x2 - 6x + 5 = 0

Answer by nyc_function(2741)

The discriminant is the name given to the expression that appears under the square root (radical) sign in the quadratic formula.
So, b^2 - 4ac is the discriminant.
Let a = -10, b = -6 and c = 5.
(-6)^2 - 4(-10)(5) =
36 + 40(5) =
36 + 200 = 236
If the discriminant is a positive number (and it is because 236 is a positive number), then there are two real solutions.
Understand?